OpenStudy (moneybaby980):

A proportional relationship between the number of pounds of cabbage (x) and the price in dollars (y) is graphed, and the ordered pair (5, 2) is on the graphed line. Part A: What is the price of 1 pound of cabbage? Show your work, including the proportion you used to determine the price. Part B: What does the ordered pair (10, 4) on the graph represent? Explain in words.

OpenStudy (anonymous):

is there a image with these problems?

OpenStudy (moneybaby980):

no

OpenStudy (anonymous):

y = k*x , where k is a constant (x,y)=(5,2) k = y/x = 2/5 so, y=(2/5)*x what is y when x is 1 pound of cabbage? y=(2/5)*1 = 2/5 (10,4) , 4 lbs of cabbage costs 10 dollars

OpenStudy (moneybaby980):

im a little confused

OpenStudy (anonymous):

the answer is 10